OpenStudy (loser66):

Let A, B \(\in M(n, \mathbb C)\)and assume B is invertible. Show that 1) \((B^{-1}AB)^k= B^{-1}A^kB\) 2) Use 1 to show \(e^{tB^{-1}AB}=B^{-1}e^{tA}B\) Please help

OpenStudy (loser66):

1) is easy. But I don't know how to do 2 @ganeshie8 @dan815

ganeshie8 (ganeshie8):

is it given that A is diagonal matrix ?

OpenStudy (loser66):

Nope, that is all.

OpenStudy (usukidoll):

invertible meaning non-singular matrix.. there's an inverse.

OpenStudy (loser66):

nvm, I got it. :) Thanks for being here.

OpenStudy (loser66):

for 1) \((B^{-1}AB)^k = (B^{-1}AB)(B^{-1}AB)........(B^{-1}AB)\) open parenthesis and \(BB^{-1}=I\) gives us the answer

OpenStudy (loser66):

for 2) by definition \(e^{tA}=\sum_{k=0}^{\infty} \dfrac{t^k}{k!}A^k\)

OpenStudy (loser66):

we replace A by \(B^{-1}AB\) then just simplify it to get the answer.

ganeshie8 (ganeshie8):

\(e^{tB^{-1}AB}=\sum_{k=0}^{\infty} \dfrac{t^k}{k!}(B^{-1}AB)^k =\sum_{k=0}^{\infty} \dfrac{t^k}{k!}B^{-1}A^kB \) ?

OpenStudy (usukidoll):

yeah.... this proof looks familiar. Miracrown helped me with one of them.

ganeshie8 (ganeshie8):

how do you conclude

OpenStudy (loser66):

@ganeshie8 sum and t, k are constants, right? hence, we can distribute inside, between \(B^{-1}\) and A

ganeshie8 (ganeshie8):

Are you saying we can take \(B^{-1}\) out ?

OpenStudy (loser66):

Yes

ganeshie8 (ganeshie8):

\(\sum_{k=0}^{\infty} \dfrac{t^k}{k!}B^{-1}A^kB =B^{-1}\left(\sum_{k=0}^{\infty} \dfrac{t^k}{k!}A^k\right)B \) ? this looks ok intuitively but idk what you're gonna use as justification

OpenStudy (loser66):

Look at the definition of \(e^{tA}\) above

ganeshie8 (ganeshie8):

Yes looked.

OpenStudy (loser66):

And we done, right?

ganeshie8 (ganeshie8):

I am asking the justification for that distribution

ganeshie8 (ganeshie8):

I think its not so obvious, at least for me it is not.

OpenStudy (dan815):